OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x + 5 through (2, 1) A. y = -3x + 7 B. y = 1/3x + 7 C. y = -1/3x + 7 D. y = 3x + 7

OpenStudy (jdoe0001):

what the slope of -> y = 1/3x + 5 ?

OpenStudy (jonnyvonny):

So, we know that for a function to be perpendicular to another, their slopes have to opposite and reciprocal ( which means that they must be the opposite sign, and raised to the -1 power [just put the slope in the denominator]).

OpenStudy (jonnyvonny):

I would love to walk this with you step by step, but i g2g...sorry.

OpenStudy (whpalmer4):

@alexwright it's time for you to respond if you want some help...

OpenStudy (anonymous):

sorry @whpalmer4 I don't know what to do on this question that's why im asking

OpenStudy (anonymous):

would it be C since it is opposite? @JonnyVonny

OpenStudy (anonymous):

is the slope (2,1) @whpalmer4

OpenStudy (whpalmer4):

No, (2,1) is the point the line goes through. \[y = \frac{1}{3}x+5\]That's slope-intercept form, from which you can "read off" the slope directly. \[y = mx+b\]\(m\) is the slope, \(b\) is the value of the y-intercept

OpenStudy (anonymous):

so which one would that be? @whpalmer4

OpenStudy (whpalmer4):

what is the slope of the line described by \[y = \frac{1}{3}x+5\]If you can't figure that out, you're not going to get anywhere with this problem (or many others). I've given you all the information you need...

OpenStudy (whpalmer4):

Compare the two equations: \[y = \frac{1}{3}x+5\]\[y=mx+b\]What is the value of \(b\) if they are equal? What is the value of \(m\) if they are equal?

OpenStudy (whpalmer4):

Having found the value of \(m\), the slope of the existing line, you can find the value of the slope of the perpendicular line: \(-1/m\) because the slopes of perpendicular lines have a product of \(-1\) (unless they are the special case of lines parallel to the x and y axes, which is not the case here). Finally, you need to use the point-slope formula to construct the equation of the line that has your new slope (which I will confusingly again call \(m\)) and passes through the point \((x_1,y_1) = (2,1)\): \[y-y_1 = m(x-x_1)\] After doing that, rearrange into the form of the answers by solving for \(y\).